# Linear combinations of latents in generative models: subspaces and beyond

### v7

@@ 5.2 Low-dimensional subspaces
$\bm{x}_{1}$ Refer to caption $\bm{x}_{2}$ Refer to caption $\bm{x}_{3}$ Refer to caption
Figure 7: Model-agnostic subspace definitions. The latents $\bm{x}_{1},\bm{x}_{2}$ and $\bm{x}_{3}$ — which via the generative model correspond to 3D designs — are converted into basis vectors and used to define a three-dimensional subspace. The left plot shows the generated designs corresponding to a ten-by-ten grid in a two-dimensional slice of the subspace, shown over a region. The right plot shows the wavedrag, as evaluated by simulation in OpenVSP (McDonald & Gloudemans, 2022) for the designs over a 100-by-100 grid in the same region, with the respective design coordinates (from the left plot) shown as black dots. We trained the SLIDE (Lyu et al., 2023) diffusion model on ShapeNet (Chang et al., 2015). LOL allows subspaces to be defined without any model-specific treatment.
In Figure 1 and Figure 6 we illustrate slices of a 5-dimensional subspace of the latent space of a flow matching model, indexing high-dimensional images of sports cars and rocking chairs, respectively. The subspaces here are defined using five images (one per desired dimension), formed using our LOL method described in Section 4 to transform the linear combinations of the corresponding projections. Our approach is dimensionality and model-agnostic, which we illustrate in Figure 7, where the same procedure is used on a completely different model without adaptations — where we define a three-dimensional subspace based on three designs in a point cloud diffusion model with mesh reconstruction. We evaluate designs in this subspace using a simple computational fluid dynamics simulation to illustrate that our approach allows objective functions to be defined over the space, in turn allowing off-the-shelf optimisation methods to be applied to search for designs. In Figure 9 in the appendix we show corresponding grids to the sports cars in Figure 1 without the proposed LOL transformation, which either leads to implausible images or the same image for each coordinate, depending on the model. In the appendix (Section G) we also include more slices and examples; including subspaces using the diffusion model Stable Diffusion 2.1 (Rombach et al., 2022).
@@ 6 Related work
Generative models with non-Gaussian priors. In Fadel et al. (2021) and  Davidson et al. (2018), in the context of normalizing flows and VAEs, respectively, Dirichlet and von Mises-Fisher prior distributions are explored with the goal of improving interpolation performance. However, these methods require training the model with the new latent distribution, which is impractical and untested for the large pretrained models we consider.
Conditional Diffusion. An additional way to control the generation process of a diffusion model is to guide the generative process with additional information, such as text and labels, to produce outputs that meet specific requirements, such as where samples are guided towards a desired class (Dhariwal & Nichol, 2021; Ho & Salimans, 2022), or to align outputs with text descriptions (Radford et al., 2021). Conditioning is complementary to latent space manipulation. For example, when making Figure 1 we used conditioning (a prompt) to constrain the generation to sports cars, however, the variation of images fulfilling this constraint is encoded in the latent space.
Low-dimensional representations. We have shown that LOL can provide expressive low-dimensional representations of latent spaces of generative models. To the best of our knowledge, the most closely related line of work was initiated in Kwon et al. (2022), where it was shown that semantic edit directions can recovered from activations of a UNet (Ronneberger et al., 2015) denoiser network during generation. Using a pretrained CLIP (Contrastive Language-Image Pre-Training) (Radford et al., 2021) model to define directions within the inner-most feature map of the UNet architecture, named h-space, they show that some semantic edits, such as adding glasses to an image of a person, correspond to linear edits in the h-space. More recently, Haas et al. (2024) demonstrate that h-space edit directions can also be found through Principle Component Analysis, and Park et al. (2023) propose a pullback metric that transfers edits in the h-space into the original latent space. However, while these three approaches demonstrate impressive editing capabilities on images, they all require a modification to the generative process and are limited to diffusion models built with UNet architectures. In our work low-dimensional representations are instead formed in closed-form as linear subspaces based on a (free) choice of latent basis vectors. These basis vectors could be obtained through various methodologies, including via the pull-back metric in Park et al. (2023), which may be interesting to explore in future work.
@@ Appendix A Optimal transport map for general latent linear combinations with independent elements
This result follows from classical optimal transport theory for univariate distributions (see Remark 2.30 in Peyré & Cuturi (2020)).
@@ Appendix B Optimal transport map for Gaussian latent linear combinations
See Remark 2.31 in Peyré & Cuturi (2020).
@@
Spherical linear interpolation (SLERP) (Shoemake, 1985) is defined as
@@
$v_{i}\in[0,1]$ and $v_{2}=1-v_{1}$ and $\cos\theta=\frac{\langle\bm{x}_{1},\bm{x}_{2}\rangle}{||\bm{x}_{1}||||\bm{x}_{2}||}$, where $\cos\theta$ is typically referred to as the cosine similarity of $\bm{x}_{1}$ and $\bm{x}_{2}$.
@@
A common case is using unit Gaussian latents (as in e.g. the models Esser et al. (2024) and Rombach et al. (2022) used in this paper), i.e. where $\bm{\mu}=\bm{0},\bm{\Sigma}=\bm{I}$. In this case it trivially follows that $\alpha\bm{\mu}=\bm{\mu}$ since $\alpha\bm{0}=\bm{0}$. We will now show that $\beta\approx 1$ in this special (i.e. unit Gaussian) case.
@@
To confirm the effectiveness of the approximations used above for values of $D$ which are typical for popular generative models that use unit Gaussian latents, we estimate the confidence interval of $\beta$ using 10k samples of $\bm{x}_{1}$ and $\bm{x}_{2}$, respectively, where $\bm{x}_{i}\sim\mathcal{N}(\bm{0},\bm{I})$. For Stable Diffusion 3 (Esser et al., 2024), a flow matching model with $D=147456$, the estimated $99\%$ confidence interval of $\beta$ is $[0.9934,1.0067]$ for $v=0.5$ where the error is largest. For Stable Diffusion 2.1 (Rombach et al., 2022), a diffusion model with $D=36864$, the corresponding confidence interval of $\beta$ is $[0.9868,1.014]$.
@@ Appendix E Normality test comparison
We compare popular alternatives for normality testing that assume a known mean and variance and that are applicable to very large sample sizes, including Kolmogorov-Smirnov (Kolmogorov, 1933), Shapiro-Wilk (Shapiro & Wilk, 1965), Chi-square (Pearson, 1900), and Cramér–von Mises (Cramér, 1928). We assess these methods on a suite of test cases of vectors which are extremely unlikely under the model’s latent distribution and which we know through failed generation violates the characteristic assumptions of the diffusion model, as well as assess the ”positive case” where the sample does follow the latent distribution, $\bm{x}_{\ast}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma})$. We report the $[0.1\%,99.9\%]$-confidence interval of the p-value produced by each respective method on $1e^{4}$ random samples in the stochastic test cases. We also include the likelihood $\mathcal{N}(\bm{x}_{\ast};\bm{\mu},\bm{\Sigma})$ and the likelihood of the norm statistic.
@@
The results are reported in Table 2. We find that Kolmogorov-Smirnov and Cramér–von Mises both succeed in reporting a lower $99.9$th percentile p-value for each tested failure case than the $0.1$th percentile assigned to real samples. Kolmogorov-Smirnov did so with the largest gap, reporting extremely low p-values for all failure cases while reporting calibrated p-values for the positive case. Shapiro-Wilk, which in the literature often is regarded as one of the most powerful tests (Razali et al., 2011), did not work well in our context in contrast to the other normality testing methods, as it produced high p-values also for several of the failure cases. An explanation may be that in our context we have sample sizes of tens of thousands (same as the dimensionality of the latent $D$), while comparisons in the literature typically focuses on smaller sample sizes, such as up to hundreds or a few thousand (Razali et al., 2011). The Chi-square test is a test for discrete distributions. However, it is commonly used on binned data (Larsen & Marx, 2005) to approximate a continuous distribution as discrete. We do this, using histograms of 30 bins. This test successfully produced very low p-values for each failure case, but also did so for many valid samples, yielding a $0.1\%$ of 0. The likelihood $\mathcal{N}(\bm{x}_{\ast};\bm{\mu},\bm{\Sigma})$ assigns high likelihoods to vectors $\bm{x}_{\ast}$ near the mode irrespective of its characteristics. We note that it fails to distinguish to failure cases from real samples, and in some cases assigns much higher likelihood to the failure cases than real data. The insufficiency of the likelihood to describe the feasibility of samples is a phenomenon present in cases of high-dimensionality and low joint dependence between these dimensions; this is due to the quickly shrinking concentration of points near the centre (or close to $\bm{\mu}$, where the likelihood is highest) as the dimensionality is increased (Talagrand, 1995; Ash, 2012; Nalisnick et al., 2019). The norm statistic we note successfully assigns low likelihoods to some failure cases, where the norm is atypical, but fails to do so in most of the tested cases.
@@
$\log\mathcal{N}(\bm{x}^{(i)};\bm{0},\bm{I})$ $[-5e^{4},-5e^{4}]$ $-3e^{4}$ $-5e^{4}$ $-3e^{4}$ $-5e^{4}$ $[-3e^{4},-3e^{4}]$ $[-5e^{4},-5e^{4}]$ $[-5e^{4},-5e^{4}]$ $[-5e^{4},-5e^{4}]$ $[-5e^{4},-5e^{4}]$
$\log\chi^{2}(||\bm{x}^{(i)}||^{2};D)$ $[-1e^{1},-7]$$\infty$- ∞ $-7$$\infty$- ∞ $-7$ $[-7e^{4},-7e^{4}]$ $[-2e^{1},-7]$ $[-7,-7]$ $[-7,-7]$ $[-7,-7]$
Shapiro-Wilk $[2e^{-4},1]$ $1$ $1$ $1$ $5e^{-72}$ $[8e^{-3},1]$ $[8e^{-3},1]$ $[1e^{-50},8e^{-19}]$ $[3e^{-105},1e^{-99}]$ $[2e^{-85},2e^{-69}]$
Chi-square $[0,1]$ 00 00 00 00 $[0,0]$ $[6e^{-149},9e^{-45}]$ $[0,1e^{-216}]$ $[0,0]$ $[0,0]$
Cramer-von-Mises $[2e^{-4},1]$ $5e^{-7}$ $0.0$ $5e^{-7}$ $1e^{-8}$ $[4e^{-7},4e^{-7}]$ $[2e^{-9},2e^{-8}]$ $[4e^{-12},7e^{-4}]$ $[2e^{-8},6e^{-8}]$ $[2e^{-11},5e^{-9}]$
Kolmogorov-Smirnov $[5e^{-4},1]$ 00 00 00 $2e^{-105}$ $[0,0]$ $[9e^{-77},9e^{-37}]$ $[1e^{-129},7e^{-10}]$ $[8e^{-252},2e^{-177}]$ $[1e^{-31},2e^{-12}]$
@@
Table 2: Normality testing of latents Reported is the $[0.1\%,99.9\%]$-confidence interval of the p-value produced by each respective normality testing method on $1e^{4}$ random samples in the stochastic test cases. We also report the log likelihood of the vector $\bm{x}^{(i)}$ under the latent distribution and the likelihood of the norm statistic. The diffusion model in Rombach et al. (2022) is used in this example. Green colour for the respective failure test case indicate the method assigning a lower p-value or likelihood to the failure case to at least 99.9% of the failure samples than the lowest 0.1% of real samples, to illustrate that it was successful in distinguishing the failure samples. Chi-square assigns low p-values to many real samples as well, which we indicate in red. The images show the generated image from a random failure sample of the test case. In (failure) Test 1 in column 1 the latent is $\bm{0}$, the mode of the latent distribution for this model $\mathcal{N}(\bm{0},\bm{I})$. In Test 2 it is the constant vector with the maximum likelihood norm of the norm sampling distribution. In Test 3 the constant vector is instead normalised to have its first two moments matching the latent’s marginal distribution. $n(\cdot)$ indicate that the vector(s) of the test case is normalised to have zero mean and unit variance. In Test 4 a linspace vector $[-1,\dots,1]^{D}$ have been similarly normalised. In Test 5 and 6 the normal samples (from the left most column) have been transformed with scaling and a bias, respectively. In Test 7 random selections of 1% of the dimensions $[1,\cdots,D]$ of the normal samples have been repeated (100 times) and subsequently normalised by $n(\cdot)$. In Test 8 and 9 a proportion of the dimensions of the normal samples have been multiplied by $5$ (with indices selected randomly) and the whole vectors are subsequently normalised by $n(\cdot)$, where the proportion is 5% and 1%, respectively.
@@ Appendix F Interpolation and centroid determination setup details
For the application of interpolation, we compare to linear interpolation (LERP), spherical linear interpolation (SLERP), and Norm-Aware Optimization (NAO) (Samuel et al., 2023), a recent approach which considers the norm of the noise vectors. In contrast to the other approaches which only involve closed-form expressions, NAO involves a numerical optimization scheme based on a discretion of a line integral. For the application of centroid determination we compare to NAO, the Euclidean centroid $\bar{\bm{x}}=\frac{1}{K}\sum_{k=1}^{K}\bm{x}_{k}$, and two transformations of the Euclidean centroid; ”standardised Euclidean”, where $\bar{\bm{x}}$ is subsequently standardised to have mean zero and unit variance, and ”mode norm Euclidean”, where $\bar{\bm{x}}$ is rescaled to have the norm equal to the (square root of the) mode of $\chi^{2}(D)$, the chi-squared distribution with $D$ degrees of freedom, which is the maximum likelihood norm given that $\bm{x}$ has been generated from a unit Gaussian with $D$ dimensions.
@@
We closely follow the evaluation protocol in Samuel et al. (2023), where we base the experiments on Stable Diffusion 2 (Rombach et al., 2022) and inversions of random images from ImageNet1k (Deng et al., 2009). We (uniformly) randomly select 50 classes, each from which we randomly select 50 unique images, and find their corresponding latents through DDIM inversion (Song et al., 2020a) using the class name as prompt. We note that the DDIM inversion can be sensitive to the number of steps. Therefore, we made sure to use a sufficient number of steps for the inversion (we used 400 steps), which we then matched for the generation; see Figure 15 for an illustration of the importance of using a sufficient number of steps. We used the guidance scale 1.0 (i.e. no guidance) for the inversion, which was then matched during generation. Note that using no guidance is important both for accurate inversion as well as to not introduce a factor (the dynamics of prompt guidance) which would be specific to the Stable Diffusion 2.1 model.
For the interpolation setup we randomly (without replacement) pair the 50 images per class into 25 pairs, forming 1250 image pairs in total. In between the ends of each respective pair, each method then is to produce three interpolation points (and images). For the NAO method, which needs additional interpolation points to approximate the line integral, we used 11 interpolation points and selected three points from these at uniform (index) distance, similar to in Samuel et al. (2023).
@@
We, as in Samuel et al. (2023), assess the methods quantitatively based on visual quality and preservation of semantics using FID distances and class prediction accuracy, respectively.
The FID distances are computed using the pytorch-fid library (Seitzer, 2020), using all evaluation images produced per method for the interpolation and centroid determination respectively, to maximize the FID distance estimation accuracy.
For the classification accuracy, we used a pre-trained classifier, the MaxViT image classification model (Tu et al., 2022) as in Samuel et al. (2023), which achieves a top-1 of 88.53% and 98.64% top-5 accuracy on the test-set of ImageNet.

### v8

@@ 5.2 Low-dimensional subspaces
$\bm{x}_{1}$ Refer to caption $\bm{x}_{2}$ Refer to caption $\bm{x}_{3}$
Figure 7: Model-agnostic subspace definitions. The latents $\bm{x}_{1},\bm{x}_{2}$ and $\bm{x}_{3}$ — which via the generative model correspond to 3D designs — are converted into basis vectors and used to define a three-dimensional subspace. The left plot shows the generated designs corresponding to a ten-by-ten grid in a two-dimensional slice of the subspace, shown over a region. The right plot shows the wavedrag, as evaluated by simulation in OpenVSP (mcdonald2022open) for the designs over a 100-by-100 grid in the same region, with the respective design coordinates (from the left plot) shown as black dots. We trained the SLIDE (lyu2023controllable) diffusion model on ShapeNet (chang2015shapenet). LOL allows subspaces to be defined without any model-specific treatment.
In Figure 1 and Figure 6 we illustrate slices of a 5-dimensional subspace of the latent space of a flow matching model, indexing high-dimensional images of sports cars and rocking chairs, respectively. The subspaces here are defined using five images (one per desired dimension), formed using our LOL method described in Section 4 to transform the linear combinations of the corresponding projections. Our approach is dimensionality and model-agnostic, which we illustrate in Figure 7, where the same procedure is used on a completely different model without adaptations — where we define a three-dimensional subspace based on three designs in a point cloud diffusion model with mesh reconstruction. We evaluate designs in this subspace using a simple computational fluid dynamics simulation to illustrate that our approach allows objective functions to be defined over the space, in turn allowing off-the-shelf optimisation methods to be applied to search for designs. In Figure 9 in the appendix we show corresponding grids to the sports cars in Figure 1 without the proposed LOL transformation, which either leads to implausible images or the same image for each coordinate, depending on the model. In the appendix (Section G) we also include more slices and examples; including subspaces using the diffusion model Stable Diffusion 2.1 (rombach2022high).
@@ 6 Related work
Generative models with non-Gaussian priors. In fadel2021principled and  davidson2018hyperspherical, in the context of normalizing flows and VAEs, respectively, Dirichlet and von Mises-Fisher prior distributions are explored with the goal of improving interpolation performance. However, these methods require training the model with the new latent distribution, which is impractical and untested for the large pretrained models we consider.
Conditional Diffusion. An additional way to control the generation process of a diffusion model is to guide the generative process with additional information, such as text and labels, to produce outputs that meet specific requirements, such as where samples are guided towards a desired class (dhariwal2021diffusion; ho2022classifier), or to align outputs with text descriptions (radford2021learning). Conditioning is complementary to latent space manipulation. For example, when making Figure 1 we used conditioning (a prompt) to constrain the generation to sports cars, however, the variation of images fulfilling this constraint is encoded in the latent space.
Low-dimensional representations. We have shown that LOL can provide expressive low-dimensional representations of latent spaces of generative models. To the best of our knowledge, the most closely related line of work was initiated in kwon2022diffusion, where it was shown that semantic edit directions can recovered from activations of a UNet (ronneberger2015u) denoiser network during generation. Using a pretrained CLIP (Contrastive Language-Image Pre-Training) (radford2021learning) model to define directions within the inner-most feature map of the UNet architecture, named h-space, they show that some semantic edits, such as adding glasses to an image of a person, correspond to linear edits in the h-space. More recently, haas2024discovering demonstrate that h-space edit directions can also be found through Principle Component Analysis, and park2023understanding propose a pullback metric that transfers edits in the h-space into the original latent space. However, while these three approaches demonstrate impressive editing capabilities on images, they all require a modification to the generative process and are limited to diffusion models built with UNet architectures. In our work low-dimensional representations are instead formed in closed-form as linear subspaces based on a (free) choice of latent basis vectors. These basis vectors could be obtained through various methodologies, including via the pull-back metric in park2023understanding, which may be interesting to explore in future work.
@@ Appendix A Optimal transport map for general latent linear combinations with independent elements
This result follows from classical optimal transport theory for univariate distributions (see Remark 2.30 in peyre2020computationaloptimaltransport).
@@ Appendix B Optimal transport map for Gaussian latent linear combinations
See Remark 2.31 in peyre2020computationaloptimaltransport.
@@
Spherical linear interpolation (SLERP) (shoemake1985animating) is defined as
@@
[…]
@@
A common case is using unit Gaussian latents (as in e.g. the models esser2024scaling and rombach2022high used in this paper), i.e. where $\bm{\mu}=\bm{0},\bm{\Sigma}=\bm{I}$. In this case it trivially follows that $\alpha\bm{\mu}=\bm{\mu}$ since $\alpha\bm{0}=\bm{0}$. We will now show that $\beta\approx 1$ in this special (i.e. unit Gaussian) case.
@@
To confirm the effectiveness of the approximations used above for values of $D$ which are typical for popular generative models that use unit Gaussian latents, we estimate the confidence interval of $\beta$ using 10k samples of $\bm{x}_{1}$ and $\bm{x}_{2}$, respectively, where $\bm{x}_{i}\sim\mathcal{N}(\bm{0},\bm{I})$. For Stable Diffusion 3 (esser2024scaling), a flow matching model with $D=147456$, the estimated $99\%$ confidence interval of $\beta$ is $[0.9934,1.0067]$ for $v=0.5$ where the error is largest. For Stable Diffusion 2.1 (rombach2022high), a diffusion model with $D=36864$, the corresponding confidence interval of $\beta$ is $[0.9868,1.014]$.
@@ Appendix E Normality test comparison
We compare popular alternatives for normality testing that assume a known mean and variance and that are applicable to very large sample sizes, including Kolmogorov-Smirnov (an1933sulla), Shapiro-Wilk (shapiro1965analysis), Chi-square (pearson1900x), and Cramér–von Mises (cramer1928composition). We assess these methods on a suite of test cases of vectors which are extremely unlikely under the model’s latent distribution and which we know through failed generation violates the characteristic assumptions of the diffusion model, as well as assess the ”positive case” where the sample does follow the latent distribution, $\bm{x}_{\ast}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma})$. We report the $0.1\%,99.9\%]-$confidence interval of the p-value produced by each respective method on $1e^{4}$ random samples in the stochastic test cases. We also include the likelihood $\mathcal{N}(\bm{x}_{\ast};\bm{\mu},\bm{\Sigma})$ and the likelihood of the norm statistic.
@@
The results are reported in Table 2. We find that Kolmogorov-Smirnov and Cramér–von Mises both succeed in reporting a lower $99.9$th percentile p-value for each tested failure case than the $0.1$th percentile assigned to real samples. Kolmogorov-Smirnov did so with the largest gap, reporting extremely low p-values for all failure cases while reporting calibrated p-values for the positive case. Shapiro-Wilk, which in the literature often is regarded as one of the most powerful tests (razali2011power), did not work well in our context in contrast to the other normality testing methods, as it produced high p-values also for several of the failure cases. An explanation may be that in our context we have sample sizes of tens of thousands (same as the dimensionality of the latent $D$), while comparisons in the literature typically focuses on smaller sample sizes, such as up to hundreds or a few thousand (razali2011power). The Chi-square test is a test for discrete distributions. However, it is commonly used on binned data (larsen2005introduction) to approximate a continuous distribution as discrete. We do this, using histograms of 30 bins. This test successfully produced very low p-values for each failure case, but also did so for many valid samples, yielding a $0.1\%$ of 0. The likelihood $\mathcal{N}(\bm{x}_{\ast};\bm{\mu},\bm{\Sigma})$ assigns high likelihoods to vectors $\bm{x}_{\ast}$ near the mode irrespective of its characteristics. We note that it fails to distinguish to failure cases from real samples, and in some cases assigns much higher likelihood to the failure cases than real data. The insufficiency of the likelihood to describe the feasibility of samples is a phenomenon present in cases of high-dimensionality and low joint dependence between these dimensions; this is due to the quickly shrinking concentration of points near the centre (or close to $\bm{\mu}$, where the likelihood is highest) as the dimensionality is increased (talagrand1995concentration; ash2012information; nalisnick2019detecting). The norm statistic we note successfully assigns low likelihoods to some failure cases, where the norm is atypical, but fails to do so in most of the tested cases.
@@
$\log\mathcal{N}(\bm{x}^{(i)};\bm{0},\bm{I})$ $5e^{4},-5e^{4}]$ $3e^{4}$ $5e^{4}$ $3e^{4}$ $5e^{4}$ $3e^{4},-3e^{4}] [$$5e^{4},-5e^{4}] [$$5e^{4},-5e^{4}] [$$5e^{4},-5e^{4}] [$$5e^{4},-5e^{4}]$
[…]
Shapiro-Wilk $2e^{-4},1]$ 11 11 11 $5e^{-72}$ $8e^{-3},1] [$$8e^{-3},1] [$$1e^{-50},8e^{-19}] [$$3e^{-105},1e^{-99}] [$$2e^{-85},2e^{-69}]$
Chi-square $[0,1]$ 0 0 0 0 $[0,0]$ $6e^{-149},9e^{-45}] [$$1e^{-216}] [0,0$][0,0] $[0,0]$
Cramer-von-Mises $2e^{-4},1]$ $5e^{-7}$ $0.0$ $5e^{-7}$ $1e^{-8}$ $4e^{-7},4e^{-7}] [$$2e^{-9},2e^{-8}] [$$4e^{-12},7e^{-4}] [$$2e^{-8},6e^{-8}] [$$2e^{-11},5e^{-9}]$
Kolmogorov-Smirnov $5e^{-4},1]$ 0 0 0 $2e^{-105}$ $[0,0]$ $9e^{-77},9e^{-37}] [$$1e^{-129},7e^{-10}] [$$8e^{-252},2e^{-177}] [$$1e^{-31},2e^{-12}]$
@@
Table 2: Normality testing of latents Reported is the $0.1\%,99.9\%]-$confidence interval of the p-value produced by each respective normality testing method on $1e^{4}$ random samples in the stochastic test cases. We also report the log likelihood of the vector $\bm{x}^{(i)}$ under the latent distribution and the likelihood of the norm statistic. The diffusion model in rombach2022high is used in this example. Green colour for the respective failure test case indicate the method assigning a lower p-value or likelihood to the failure case to at least 99.9% of the failure samples than the lowest 0.1% of real samples, to illustrate that it was successful in distinguishing the failure samples. Chi-square assigns low p-values to many real samples as well, which we indicate in red. The images show the generated image from a random failure sample of the test case. In (failure) Test 1 in column 1 the latent is $\bm{0}$, the mode of the latent distribution for this model $\mathcal{N}(\bm{0},\bm{I})$. In Test 2 it is the constant vector with the maximum likelihood norm of the norm sampling distribution. In Test 3 the constant vector is instead normalised to have its first two moments matching the latent’s marginal distribution. $n(\cdot)$ indicate that the vector(s) of the test case is normalised to have zero mean and unit variance. In Test 4 a linspace vector $1,\dots,1]^{D}$ have been similarly normalised. In Test 5 and 6 the normal samples (from the left most column) have been transformed with scaling and a bias, respectively. In Test 7 random selections of 1% of the dimensions $1,\cdots,D]$ of the normal samples have been repeated (100 times) and subsequently normalised by $n(\cdot)$. In Test 8 and 9 a proportion of the dimensions of the normal samples have been multiplied by 55 (with indices selected randomly) and the whole vectors are subsequently normalised by $n(\cdot)$, where the proportion is 5% and 1%, respectively.
@@ Appendix F Interpolation and centroid determination setup details
For the application of interpolation, we compare to linear interpolation (LERP), spherical linear interpolation (SLERP), and Norm-Aware Optimization (NAO) (samuel2024norm), a recent approach which considers the norm of the noise vectors. In contrast to the other approaches which only involve closed-form expressions, NAO involves a numerical optimization scheme based on a discretion of a line integral. For the application of centroid determination we compare to NAO, the Euclidean centroid $\bar{\bm{x}}=\frac{1}{K}\sum_{k=1}^{K}\bm{x}_{k}$, and two transformations of the Euclidean centroid; ”standardised Euclidean”, where $\bar{\bm{x}}$ is subsequently standardised to have mean zero and unit variance, and ”mode norm Euclidean”, where $\bar{\bm{x}}$ is rescaled to have the norm equal to the (square root of the) mode of $\chi^{2}(D)$, the chi-squared distribution with $D$ degrees of freedom, which is the maximum likelihood norm given that $\bm{x}$ has been generated from a unit Gaussian with $D$ dimensions.
@@
We closely follow the evaluation protocol in samuel2024norm, where we base the experiments on Stable Diffusion 2 (rombach2022high) and inversions of random images from ImageNet1k (deng2009imagenet). We (uniformly) randomly select 50 classes, each from which we randomly select 50 unique images, and find their corresponding latents through DDIM inversion (song2020denoising) using the class name as prompt. We note that the DDIM inversion can be sensitive to the number of steps. Therefore, we made sure to use a sufficient number of steps for the inversion (we used 400 steps), which we then matched for the generation; see Figure 15 for an illustration of the importance of using a sufficient number of steps. We used the guidance scale 1.0 (i.e. no guidance) for the inversion, which was then matched during generation. Note that using no guidance is important both for accurate inversion as well as to not introduce a factor (the dynamics of prompt guidance) which would be specific to the Stable Diffusion 2.1 model.
For the interpolation setup we randomly (without replacement) pair the 50 images per class into 25 pairs, forming 1250 image pairs in total. In between the ends of each respective pair, each method then is to produce three interpolation points (and images). For the NAO method, which needs additional interpolation points to approximate the line integral, we used 11 interpolation points and selected three points from these at uniform (index) distance, similar to in samuel2024norm.
@@
We, as in samuel2024norm, assess the methods quantitatively based on visual quality and preservation of semantics using FID distances and class prediction accuracy, respectively.
The FID distances are computed using the pytorch-fid library (Seitzer2020FID), using all evaluation images produced per method for the interpolation and centroid determination respectively, to maximize the FID distance estimation accuracy.
For the classification accuracy, we used a pre-trained classifier, the MaxViT image classification model (tu2022maxvit) as in samuel2024norm, which achieves a top-1 of 88.53% and 98.64% top-5 accuracy on the test-set of ImageNet.